# Cohomology of Lie algebroids over Algebraic spaces

### v6

@@ 4. Logarithmic de Rham cohomology
We compute algebraic (analytic) de Rham cohomology groups of some family of non-singular varieties $Y_{t}$ for $t\in\mathbb{C}\setminus\{0\}$ and its associated singular variety $Y:=Y_{0}$ (which appears as principal divisors [MS22]), i.e. cohomology of the Lie algebroid $\mathcal{T}_{Y_{t}}$. Then, we compute hypercohomology of the logarithmic de Rham complexes [CM99], which is a simplified version of the Lie algebroid cohomology for $\mathcal{T}_{X}(-logY)$ [Pym13], since we work over some free divisors $Y\subset X$ [CJNMM96, Pym13].
@@
We can use algebraic de Rham theorem [Dan89, Ste15] to get algebraic de Rham cohomology of $Y$ by the singular cohomology of $S^{1}$. Now we recall the computation which helps us to follow the associated singular case when $t=0$ (which is a normal crossing divisor).
@@
This is studied in [Hub16], as a topological invariant for non singular spaces, but the analogous computation for the associated singular case is not derived there.
Note that the affine algebraic set $Y$ can be considered as an analytic space. Moreover, this space can be viewed as a principal divisor [MS22, Pym13].
We consider the associated logarithmic de Rham cohomology [CJNMM96, CM99] in the context of complex geometry.
@@
Here, $\mathcal{O}_{X}$ is the sheaf of holomorphic functions on $X:=\mathbb{C}^{2}$ and $(Y:=V(\mathcal{I}),\mathcal{O}_{Y}:=\mathcal{O}_{X}/{\mathcal{I}})$ is the analytic space associated with the principal ideal sheaf $\mathcal{I}=\langle xy-1\rangle$ of $\mathcal{O}_{X}$. The sheaf of logarithmic derivations for $Y$ in $X$ is $\mathcal{T}_{X}(-logY)=\langle\{x\partial_{x}-y\partial_{y},f~{}\nabla f\}\rangle$ where $\nabla f=y\partial_{x}+x\partial_{y}$ (since $f=xy-1)$. The sheaf of logarithmic $1$-forms for $Y$ in $X$ (i.e. meromorphic $1$-forms $\omega$ of $X$ with poles along the divisor $Y$ such that $f\omega\in\Omega^{1}_{X}$ and $df\wedge\omega\in\Omega^{2}_{X}$ [CM99]) is
@@
To determine hypercohomology [Dan89, Ste15] of the logarithmic de Rham complex
@@
Thus, we can view the normal crossing divisor $Y_{0}:=\{(x,y)\in\mathbb{C}^{2}\mid xy=0\}$ (a singular analytic space and a free divisor) as deformation of family of rectangular hyperbolas $Y_{t}$ (appears in [Pfl96] as deformation of a scheme).
@@
Note that $Y_{0}$ is of same homotopy type to a point, thus singular cohomologies of $Y_{0}$ vanishes in all higher dimensions ($\geq 1$). Fortunately, for this singular variety the algebraic de Rham cohomology and the singular cohomology are same (see [Gil12]).
This approach does not provide a topological invariant for all singular algebraic varieties, to resolve this problem a different approach is considered in [Hub16].
We consider the associated cohomology in complex geometry context (see [MS22]).
@@
Note that, the sheaf $\mathcal{T}_{X}(-logY_{0})$ has a canonical Lie algebroid structure, and $Y_{0}$ is a free divisor [Pym13, MS22]. Since, $log$ $x$ and $log$ $y$ both are not in $\mathcal{O}_{X}$, thus $\frac{dx}{x}$ and $\frac{dy}{y}$ are not in $\mathscr{I}m$ $d^{0}$ but both are in $\mathscr{K}er$ $d^{1}$. Similarly, $\frac{dx}{x}\wedge\frac{dy}{y}$ is in $\mathscr{K}er$ $d^{2}$ but not in $\mathscr{I}m$ $d^{1}$. Other choices for cocycles are appears as coboundaries.
@@
Now, consider the holomorphic Lie algebroid cohomology [BMRT15]
@@
(using the standard Lemma of Atiyah-Hodge [Ste15]), similar results appear in [CJNMM96].
@@
In general for smooth algebraic variety and complex manifold, to compute its cohomology we need to consider sheaf of de Rham complex because the space of global sections not necessarily captures the whole information. In these cases we can use Čech cohomology [Dan89, Ram05] by considering an affine open cover or Stein open cover ((((known as good open cover)))) [BMRT15, Ste15] accordingly.
@@ 5. Hochschild Cohomology for a Lie algebroid over $a$-spaces
In this section, we define Hochschild cohomology of an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra (see Section 2.3), using (Hopf-)Hochschild cohomology of a left bialgebroid [KP11] as local descriptions. In particular, we describe the cases associated with universal enveloping algebroid and jet algebroid of a locally free Lie algebroid over an $a$-space. The cohomology groups can be computed using suitable standard complexes. We prove a version of Hochschild-Kostant-Rosenberg (HKR) theorem and dual HKR theorem. It is done by following the local counterpart as can be found in [KP11], where the authors proved (dual) HKR theorem for (finitely generated) projective Lie-Rinehart algebras. Moreover, we present the HKR theorem for sheaf of logarithmic differential operators and sheaf of noncommuative differential operators.
@@ 5.1. Hochschild hypercohomology of an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra
where $C^{\bullet}(\mathcal{A}(U))=((\mathcal{A}(U))^{\otimes_{\mathcal{O}_{X}(U)}^{\bullet}},b)$ is the Hochschild cochain complex of the $\mathcal{O}_{X}(U)/{\mathbb{K}}$-bialgebra $\mathcal{A}(U)$ with the differential $b$ defined by
$b(a_{1}\otimes\cdots\otimes a_{n})=1\otimes a_{1}\otimes\cdots\otimes a_{n}+\sum^{n}_{i=1}(-1)^{i}a_{1}\otimes\cdots\otimes\Delta(a_{i})\otimes\cdots\otimes a_{n}$
$+(-1)^{n+1}a_{1}\otimes\cdots\otimes a_{n}\otimes 1.$
@@
where $Cob^{\bullet}({\mathcal{A}(U)}):=((\mathcal{A}(U))^{\otimes_{\mathcal{O}_{X}(U)}^{\bullet+1}},b^{\prime})$ is the cobar complex for $\mathcal{O}_{X}(U)$ in the category of left $\mathcal{A}(U)$-comodules with the differential $b^{\prime}$ defined by
@@
call it cobar resolution of $\mathcal{A}$. Then by applying cotensor product functor $\mathcal{O}_{X}\Box_{\mathcal{A}}-$ (sheafifying the cotensor functor described in [KP11]) on it and considering hypercohomology (i.e. applying hypercohomology $\mathbb{H}^{\bullet}(X,-)$ functor on the induced cochain complex $\mathcal{O}_{X}\Box_{\mathcal{A}}\mathscr{C}ob^{\bullet}(\mathcal{A})$ of $\mathbb{K}_{X}$-vector spaces), we get the following isomorphism (sheafifying the local descriptions of relationship among Hochschild cochain complex and cobar resolution [KP11] and considering their associated hypercohomologies)
@@
Now, we describe Hochschild cohomology for some special $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras, namely for the universal enveloping algebroid $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ [MS22] and the jet algebroid $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ [MS22] of a (locally free) Lie algebroid $\mathcal{L}$ over a special $a$-space (as mentioned in Note 2.2) $(X,\mathcal{O}_{X})$.
@@ 5.2. Hochschild hypercohomology of universal enveloping algebroid
The sheaf of $\mathcal{L}$-poly-vector fields on $X$ is defined as $\mathcal{T}^{poly}_{\mathcal{L}}(\mathcal{O}_{X}):=\bigoplus_{i}\wedge^{i}\mathcal{L}$. This generalizes the space of multisections of the tangent bundle (i.e. multi-vector fields) descried for different $a$-spaces occurs in geometries. It has a canonical sheaf of Gerstenhaber algebra structure on $X$. For the case of a smooth manifold $X$, Kontsevich introduced the (sheaf of) poly-differential operators on $X$. It induces a subcomplex of the Hochschild cochain complex for $\mathcal{O}_{X}$. Its analogue in the context of Lie algebroid is the sheaf of $\mathcal{L}$-poly differential operators $\mathcal{D}^{poly}_{\mathcal{L}}(\mathcal{O}_{X}):=\bigoplus_{i}\mathscr{U}(\mathcal{O}_{X},\mathcal{L})^{\otimes_{\mathcal{O}_{X}}^{i}}$. It has a canonical sheaf of Gerstenhaber algebra structure on $X$. In particular, when $\mathcal{L}=\mathcal{T}_{X}$ we denote these sheaves by the standard notions $\mathcal{T}^{poly}_{X}$ and $\mathcal{D}^{poly}_{X}$ respectively. The so-called HKR map is a quasi-isomorphism between $\mathcal{T}^{poly}_{X}$ and $\mathcal{D}^{poly}_{X}$ (of differential graded $\mathbb{K}_{X}$-vector spaces). See [CVdB10] for more generalities.
Using ideas from proof of HKR theorem (and canonical PBW coalgebra isomorphism) for Lie-Rinehart algebras [KP11] we get the following results. To state these results first we need to recall some notations associated with a $(\mathbb{K},R)$-Lie-Rinehart algebra $L$ and an $R/{\mathbb{K}}$-bialgebra $A$. The symmetric algebra of $L$ over $R$ (using underlying $R$-module structure) is denoted by $S_{R}L$ and $S_{R}L^{*}$ is the symmetric algbra for the $R$-module $L^{*}:=Hom_{R}(L,R)$, the universal enveloping algebra of the $(\mathbb{K},R)$-Lie-Rinehart algebra $L$ is denoted by $\mathcal{U}(R,L)$ and its dual is the jet algebra $\mathcal{J}(R,L)$, the cobar resolution of $A$ (using underlying $R$-coalgebra structure) is denoted by $Cob^{\bullet}(A)$ and the associated cohomology (applying the functor $R$ $\Box_{A}$ using cotensor product) is Hochschild cohomology of the $R/{\mathbb{K}}$-bialgebra $A$ denoted by $HH^{\bullet}(A)$ (in our cases $A$ is either $S_{R}L$, $\mathcal{U}(R,L)$ or $S_{R}L^{*}$, $\mathcal{J}(R,L)$ with canonical $R/\mathbb{K}$-bialgebra structure [MM10, CVdB10, CRvdB10, KP11]).
We know that the sheaf of symmetric algebras $\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}$ can be viewed as universal enveloping algebroid of the Lie algebroid $\mathcal{L}$ if the $\mathcal{O}_{X}$-module $\mathcal{L}$ is equipped with zero bracket and zero anchor (on each space of sections) [MS22]. Recall that we have PBW presheaf homomorphism of presheaf of $\mathcal{O}_{X}$-algebas given as
@@
where $f$ is a section of $\mathcal{O}_{X}$, $D_{i}$ is a section of $\mathcal{L}$ and $\bar{D}_{i}$ is its associated image in $\mathcal{U}(\mathcal{O}_{X},\mathcal{L})$ $(i=1,2,\dots,k)$. Sheafification of this map gives the generalized PBW map between the associated sheaves (without considering the graded quotient) [MS22]. In addition, if $\mathcal{L}$ is locally free $\mathcal{O}_{X}$-module then we get following results.
@@
The $(\mathbb{K},\mathcal{O}_{X}(U_{x}))$-Lie-Rinehart algebra $\mathcal{L}(U_{x})$ is a free $\mathcal{O}_{X}(U_{x})$-module for each special open set $U_{x}$. As in[KP11], we consider the isomorphism of cochain complexes associated with each special open set $U_{x}$ given as $Cob^{\bullet}(\mathcal{U}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x}))\overset{\cong}{\rightarrow}Cob^{\bullet}(S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})).$ The isomorphism induces an isomorphism of graded vector spaces
@@
Here we apply the notion of hypercohomology (Note 3.2) for the cochain complex of sheaves $\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{F}$. For each sheaf of ${\mathcal{O}_{X}}$-modules (or $\mathcal{O}_{X}$-module) $\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F}$, we have a flabby Godement resolution (given by a quasi-isomorphism)
@@
For the complex of sheaves $\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{F}$, we consider the corresponding bicomplex of sheaves $\mathscr{C}^{\bullet}(\mathcal{F}):=(\mathscr{C}^{j}(\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F}))_{i,j\geq 0}$. The original complex can be embedded in the total complex $\mathscr{K}^{\bullet}:=tot(\mathscr{C}^{\bullet}(\mathcal{F}))$. Moreover, this embedding is a quasi-isomorphism. The cohomology of the associated complex of global sections $\mathscr{K}^{\bullet}(X)=tot(\mathscr{C}^{\bullet}(\mathcal{F}))(X)$ is the hypercohomology of the complex $\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{F}$. Since here differential of the complex is zero, thus the computation of cohomology of the total complex is become a simpler one. It is directly expressed through the sheaf cohomologies $H^{j}(X,\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F})$ as stated in the Lemma.
@@
(Generalized HKR theorem) Let $\mathcal{L}$ be a locally free Lie algebroid over $(X,\mathcal{O}_{X})$. Then the anti-symmetrization map (known as HKR morphism)
@@
Assume first $\mathcal{L}$ to be a locally free Lie algebroid of finite rank. Since for each $x\in X$ we have an open set $U_{x}$ containing $x$ with $\mathcal{L}|_{U_{x}}$ is a free $\mathcal{O}_{X}|_{U_{x}}$-module, thus on each $U_{x}$ we use results from the proof of the HKR Theorem [KP11] for the $(\mathbb{K},\mathcal{O}_{X}(U_{x}))$-Lie-Rinehart algebra $\mathcal{L}(U_{x})$. Then, for each special open sets $U_{x}$ we have isomorphisms of cochain complexes of left $S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})$-comodules
@@
compatible with restrictions, where $\tilde{K}^{\bullet}_{U}\mathcal{L}:=S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)\otimes_{\mathcal{O}_{X}(U)}\wedge^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U)$ and the differential $\partial_{U}:\tilde{K}^{\bullet}_{U}\mathcal{L}\rightarrow\tilde{K}^{\bullet+1}_{U}\mathcal{L}$ is described in [KP11], for any open set $U$ in $X$. Indeed, the dual Koszul-Rinehart complex for $\mathcal{L}(U)$ is given as $\tilde{K}^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U):=(\tilde{K}^{\bullet}_{U}\mathcal{L},$ $\partial_{U})$.
@@
for each special open set $U_{x}$ (using results from [KP11]). Since both $\mathbb{H}H^{\bullet}$ (which is isomorphic to the derived functor $Cotor^{\bullet}$) and the functor $\mathcal{S}$ commute with inductive limits over a filtered ordered set, we get the result using the result follows from the locally free $\mathcal{O}_{X}$-module of finite rank case (local counterpart is described in [KP11]). ∎
@@
The HKR morphism is not a ring isomorphism, but composing the HKR-morphism together with the Todd genus provides a ring isomorphism (moreover it provides a canonical Gerstenhaber algebra isomorphism). It is required in the study of formality (or quantization) for Lie algebroids [CVdB10].

### v7

@@ 4. Logarithmic de Rham cohomology
We compute algebraic (analytic) de Rham cohomology groups of a family of non-singular varieties $Y_{t}$ for $t\in\mathbb{C}\setminus\{0\}$ and their associated singular variety $Y:=Y_{0}$ (which appears as principal divisors [29]), i.e. cohomology of the Lie algebroid $\mathcal{T}_{Y_{t}}$ for $t\in\mathbb{C}$. Then, we compute hypercohomology of the logarithmic de Rham complexes [27], which simplifies to the Lie algebroid cohomology for $\mathcal{T}_{X}(-logY)$ [32], as we work over some free divisors $Y\subset X$ [4, 32].
@@
We can use algebraic de Rham theorem [8, 36] to get algebraic de Rham cohomology of $Y$ by the singular cohomology of $S^{1}$. Now we recall the computation which helps us to follow the associated singular case when $t=0$ (which is a normal crossing divisor).
@@
This is studied in [15], as a topological invariant for non singular spaces, but the analogous computation for the associated singular case is not derived there.
Note that the affine algebraic set $Y$ can be considered as an analytic space. Moreover, this space can be viewed as a principal divisor [29, 32].
We consider the associated logarithmic de Rham cohomology [4, 27] in the context of complex geometry.
@@
Here, $\mathcal{O}_{X}$ is the sheaf of holomorphic functions on $X:=\mathbb{C}^{2}$ and $(Y:=V(\mathcal{I}),\mathcal{O}_{Y}:=\mathcal{O}_{X}/{\mathcal{I}})$ is the analytic space associated with the principal ideal sheaf $\mathcal{I}=\langle xy-1\rangle$ of $\mathcal{O}_{X}$. The sheaf of logarithmic derivations for $Y$ in $X$ is $\mathcal{T}_{X}(-logY)=\langle\{x\partial_{x}-y\partial_{y},f~{}\nabla f\}\rangle$ where $\nabla f=y\partial_{x}+x\partial_{y}$ ((((since $f=xy-1)$. The sheaf of logarithmic $1$-forms for $Y$ in $X$ (i.e. meromorphic $1$-forms $\omega$ of $X$ with poles along the divisor $Y$ such that $f\omega\in\Omega^{1}_{X}$ and $df\wedge\omega\in\Omega^{2}_{X}$ [27]) is
@@
To determine hypercohomology [8, 36] of the logarithmic de Rham complex
@@
Thus, we can view the normal crossing divisor $Y_{0}:=\{(x,y)\in\mathbb{C}^{2}\mid xy=0\}$ (a singular analytic space and a free divisor) as deformation of family of rectangular hyperbolas $Y_{t}$ (appears in [31] as deformation of a scheme).
@@
Note that $Y_{0}$ is of same homotopy type to a point, thus singular cohomologies of $Y_{0}$ vanishes in all higher dimensions ($\geq 1$). Fortunately, for this singular variety the algebraic de Rham cohomology and the singular cohomology are same (see [12]).
This approach does not provide a topological invariant for all singular algebraic varieties, to resolve this problem a different approach is considered in [15].
We consider the associated cohomology in complex geometry context (see [29]).
@@
Note that, the sheaf $\mathcal{T}_{X}(-logY_{0})$ has a canonical Lie algebroid structure, and $Y_{0}$ is a free divisor [32, 29]. Since, $log$ $x$ and $log$ $y$ both are not in $\mathcal{O}_{X}$, thus $\frac{dx}{x}$ and $\frac{dy}{y}$ are not in $\mathscr{I}m$ $d^{0}$ but both are in $\mathscr{K}er$ $d^{1}$. Similarly, $\frac{dx}{x}\wedge\frac{dy}{y}$ is in $\mathscr{K}er$ $d^{2}$ but not in $\mathscr{I}m$ $d^{1}$. Other choices for cocycles are appears as coboundaries.
@@
Now, consider the holomorphic Lie algebroid cohomology [1]
@@
(using the standard Lemma of Atiyah-Hodge [36]), similar results appear in [4].
@@
In general for smooth algebraic variety and complex manifold, to compute its cohomology we need to consider sheaf of de Rham complex because the space of global sections not necessarily captures the whole information. In these cases we can use Čech cohomology [8, 33] by considering an affine open cover or Stein open cover ((((known as good open cover)))) [1, 36] accordingly.
@@ 5. Hochschild Cohomology for a Lie algebroid over $a$-spaces
In this section, we define Hochschild cohomology of an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra (see Section 2.3), using (Hopf-)Hochschild cohomology of a left bialgebroid [19] as local descriptions. In particular, we describe the cases associated with universal enveloping algebroid and jet algebroid of a locally free Lie algebroid over an $a$-space. The cohomology groups can be computed using suitable standard complexes. We prove a version of Hochschild-Kostant-Rosenberg (HKR) theorem and dual HKR theorem. It is done by following the local counterpart as can be found in [19], where the authors proved (dual) HKR theorem for (finitely generated) projective Lie-Rinehart algebras. Moreover, we present the HKR theorem for the sheaf of logarithmic differential operators and the sheaf of noncommuative differential operators.
@@ 5.1. Hochschild hypercohomology of an $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra
where $C^{\bullet}(\mathcal{A}(U))=((\mathcal{A}(U))^{\otimes_{\mathcal{O}_{X}(U)}^{\bullet}},b)$ is the Hochschild cochain complex of the $\mathcal{O}_{X}(U)/{\mathbb{K}}$-bialgebra $\mathcal{A}(U)$ with the differential $b$ defined for any $a_{1},\dots,a_{n}\in\mathcal{A}(U)$ as follows:
$b(a_{1}\otimes\cdots\otimes a_{n})=1\otimes a_{1}\otimes\cdots\otimes a_{n}+\sum^{n}_{i=1}(-1)^{i}a_{1}\otimes\cdots\otimes\Delta(a_{i})\otimes\cdots\otimes a_{n}$ $+(-1)^{n+1}a_{1}\otimes\cdots\otimes a_{n}\otimes 1.$
@@
where $Cob^{\bullet}({\mathcal{A}(U)}):=((\mathcal{A}(U))^{\otimes_{\mathcal{O}_{X}(U)}^{\bullet+1}},b^{\prime})$ is the cobar complex for $\mathcal{O}_{X}(U)$ in the category of left $\mathcal{A}(U)$-comodules with the differential $b^{\prime}$ defined for any $a_{0},\dots,a_{n}\in\mathcal{A}(U)$ as follows:
@@
call it cobar resolution of $\mathcal{A}$. Then by applying cotensor product functor $\mathcal{O}_{X}\Box_{\mathcal{A}}-$ (sheafifying the cotensor functor described in [19]) on it and considering hypercohomology (i.e. applying hypercohomology $\mathbb{H}^{\bullet}(X,-)$ functor on the induced cochain complex $\mathcal{O}_{X}\Box_{\mathcal{A}}\mathscr{C}ob^{\bullet}(\mathcal{A})$ of $\mathbb{K}_{X}$-vector spaces), we get the following isomorphism (sheafifying the local descriptions of relationship among Hochschild cochain complex and cobar resolution [19] and considering their associated hypercohomologies)
@@
Now, we describe Hochschild cohomology for some special $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras, namely for the universal enveloping algebroid $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ [29] and the jet algebroid $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ [29] of a (locally free) Lie algebroid $\mathcal{L}$ over a special $a$-space (as mentioned in Note 2.2) $(X,\mathcal{O}_{X})$.
@@ 5.2. Hochschild hypercohomology of universal enveloping algebroid
The sheaf of $\mathcal{L}$-poly-vector fields on $X$ is defined as $\mathcal{T}^{poly}_{\mathcal{L}}(\mathcal{O}_{X}):=\bigoplus_{i}\wedge^{i}\mathcal{L}$. This generalizes the space of multisections of the tangent bundle (i.e. multi-vector fields) descried for different $a$-spaces occurs in geometries. It has a canonical sheaf of Gerstenhaber algebra structure on $X$. For the case of a smooth manifold $X$, Kontsevich introduced the (sheaf of) poly-differential operators on $X$. It induces a subcomplex of the Hochschild cochain complex for $\mathcal{O}_{X}$. Its analogue in the context of Lie algebroid is the sheaf of $\mathcal{L}$-poly differential operators $\mathcal{D}^{poly}_{\mathcal{L}}(\mathcal{O}_{X}):=\bigoplus_{i}\mathscr{U}(\mathcal{O}_{X},\mathcal{L})^{\otimes_{\mathcal{O}_{X}}^{i}}$. It has a canonical sheaf of Gerstenhaber algebra structure on $X$. In particular, when $\mathcal{L}=\mathcal{T}_{X}$ we denote these sheaves by the standard notions $\mathcal{T}^{poly}_{X}$ and $\mathcal{D}^{poly}_{X}$ respectively. The so-called HKR map is a quasi-isomorphism between $\mathcal{T}^{poly}_{X}$ and $\mathcal{D}^{poly}_{X}$ (of differential graded $\mathbb{K}_{X}$-vector spaces). See [7] for more generalities.
Using ideas from proof of HKR theorem (and canonical PBW coalgebra isomorphism) for Lie-Rinehart algebras [19] we get the following results. To state these results first we need to recall some notations associated with a $(\mathbb{K},R)$-Lie-Rinehart algebra $L$ and an $R/{\mathbb{K}}$-bialgebra $A$. The symmetric algebra of $L$ over $R$ (using underlying $R$-module structure) is denoted by $S_{R}L$ and $S_{R}L^{*}$ is the symmetric algbra for the $R$-module $L^{*}:=Hom_{R}(L,R)$, the universal enveloping algebra of the $(\mathbb{K},R)$-Lie-Rinehart algebra $L$ is denoted by $\mathcal{U}(R,L)$ and its dual is the jet algebra $\mathcal{J}(R,L)$, the cobar resolution of $A$ (using underlying $R$-coalgebra structure) is denoted by $Cob^{\bullet}(A)$ and the associated cohomology (applying the functor $R$ $\Box_{A}$ using cotensor product) is Hochschild cohomology of the $R/{\mathbb{K}}$-bialgebra $A$ denoted by $HH^{\bullet}(A)$ (in our cases $A$ is either $S_{R}L$, $\mathcal{U}(R,L)$ or $S_{R}L^{*}$, $\mathcal{J}(R,L)$ with canonical $R/\mathbb{K}$-bialgebra structure [26, 7, 5, 19]).
We know that the sheaf of symmetric algebras $\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L}$ can be viewed as universal enveloping algebroid of the Lie algebroid $\mathcal{L}$ if the $\mathcal{O}_{X}$-module $\mathcal{L}$ is equipped with zero bracket and zero anchor (on each space of sections) [29]. Recall that we have PBW presheaf homomorphism of presheaf of $\mathcal{O}_{X}$-algebas given as
@@
where $f$ is a section of $\mathcal{O}_{X}$, $D_{i}$ is a section of $\mathcal{L}$ and $\bar{D}_{i}$ is its associated image in $\mathcal{U}(\mathcal{O}_{X},\mathcal{L})$ $(i=1,2,\dots,k)$. Sheafification of this map gives the generalized PBW map between the associated sheaves (without considering the graded quotient) [29]. In addition, if $\mathcal{L}$ is locally free $\mathcal{O}_{X}$-module then we get following results.
@@
The $(\mathbb{K},\mathcal{O}_{X}(U_{x}))$-Lie-Rinehart algebra $\mathcal{L}(U_{x})$ is a free $\mathcal{O}_{X}(U_{x})$-module for each special open set $U_{x}$. As in[19], we consider the isomorphism of cochain complexes associated with each special open set $U_{x}$ given as $Cob^{\bullet}(\mathcal{U}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x}))\overset{\cong}{\rightarrow}Cob^{\bullet}(S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})).$ The isomorphism induces an isomorphism of graded vector spaces
@@
Here we apply the notion of hypercohomology (Note 3.2) for the cochain complex of sheaves $\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{F}$. For each sheaf of ${\mathcal{O}_{X}}$-modules (or $\mathcal{O}_{X}$-module) $\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F}$, we have an injective resolution (given by the flabby Godement resolution) $\mathscr{C}^{\bullet}(\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F})$, provides a quasi-isomorphism
@@
The corresponding bicomplex for the complex of sheaves $\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{F}$ is $\mathscr{C}^{\bullet}(\mathcal{F}):=(\mathscr{C}^{j}(\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F}))_{i,j\geq 0}$. The original complex can be embedded in the total complex $\mathscr{K}^{\bullet}:=tot(\mathscr{C}^{\bullet}(\mathcal{F}))$. Moreover, this embedding is a quasi-isomorphism. The cohomology of the associated complex of global sections $\mathscr{K}^{\bullet}(X)=tot(\mathscr{C}^{\bullet}(\mathcal{F}))(X)$ is the hypercohomology of the complex $\wedge^{\bullet}_{\mathcal{O}_{X}}\mathcal{F}$. Since here differential of the complex is zero, thus the computation of cohomology of the total complex is become a simpler one. It is directly expressed through the sheaf cohomologies $H^{j}(X,\wedge^{i}_{\mathcal{O}_{X}}\mathcal{F})$ as stated in the Lemma.
@@
((((Generalized HKR theorem)))) Let $\mathcal{L}$ be a locally free Lie algebroid over $(X,\mathcal{O}_{X})$. Then the anti-symmetrization map (known as HKR morphism)
@@
Assume first $\mathcal{L}$ to be a locally free Lie algebroid of finite rank. Since for each $x\in X$ we have an open set $U_{x}$ containing $x$ with $\mathcal{L}|_{U_{x}}$ is a free $\mathcal{O}_{X}|_{U_{x}}$-module, thus on each $U_{x}$ we use results from the proof of the HKR Theorem [19] for the $(\mathbb{K},\mathcal{O}_{X}(U_{x}))$-Lie-Rinehart algebra $\mathcal{L}(U_{x})$. Then, for each special open sets $U_{x}$ we have isomorphisms of cochain complexes of left $S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})$-comodules
@@
compatible with restrictions, where $\tilde{K}^{\bullet}_{U}\mathcal{L}:=S_{\mathcal{O}_{X}(U)}\mathcal{L}(U)\otimes_{\mathcal{O}_{X}(U)}\wedge^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U)$ and the differential $\partial_{U}:\tilde{K}^{\bullet}_{U}\mathcal{L}\rightarrow\tilde{K}^{\bullet+1}_{U}\mathcal{L}$ is described in [19], for any open set $U$ in $X$. Indeed, the dual Koszul-Rinehart complex for $\mathcal{L}(U)$ is given as $\tilde{K}^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U):=(\tilde{K}^{\bullet}_{U}\mathcal{L},$ $\partial_{U})$.
@@
for each special open set $U_{x}$ (using results from [19]). Since both $\mathbb{H}H^{\bullet}$ (which is isomorphic to the derived functor $Cotor^{\bullet}$) and the functor $\mathcal{S}$ commute with inductive limits over a filtered ordered set, we get the result using the result follows from the locally free $\mathcal{O}_{X}$-module of finite rank case (local counterpart is described in [19]). ∎
@@
The HKR morphism is not a ring isomorphism, but composing the HKR-morphism together with the Todd genus provides a ring isomorphism (moreover it provides a canonical Gerstenhaber algebra isomorphism). It is required in the study of formality (or quantization) for Lie algebroids [7].